OpenStudy (solomonzelman):

special tricks? ~ cyclic integrals? ~ cube in latex? ~ very big triangle and circle in latex?

OpenStudy (solomonzelman):

I made up a trick for a \(\normalsize\color{royalblue}{ \rm square }\) . \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\Huge\frac{~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}}\) but how about the afore-listed things?

OpenStudy (aripotta):

\(\Huge \triangle \bigcirc\) lol idk

OpenStudy (solomonzelman):

that is `\triangle`, `\circ` with a size `\Huge`, but I need as big as this square I posted.

OpenStudy (solomonzelman):

Ari Potta, and do you by chance know about cyclic integrals?

OpenStudy (aripotta):

no i don't :/ sorry i can't help lol. i'm just starting out in latex

OpenStudy (unklerhaukus):

`\[ \int, \oint, \iint, \iiint \]` \[ \int, \oint, \iint, \iiint \]

OpenStudy (solomonzelman):

http://en.wikipedia.org/wiki/Integral_symbol tnx

OpenStudy (confluxepic):



sammixboo (sammixboo):

I've looked all over for a LaTeX Cube, but I can't find it. Anyways, here is a dagger \(\huge \dagger\)

OpenStudy (solomonzelman):

I ain't need a dagger, esp considering the fact that I am Jewish, and IT really looks like a cross.

OpenStudy (solomonzelman):

tnx though

OpenStudy (confluxepic):

@sammixboo You found my dagger. Didn't you. :) lol

OpenStudy (anonymous):

Haha, nah :P I found it while playing around with a whole bunch of LaTeX